OpenStudy (anonymous):

If a spring stores 33.5 J of energy when it's compressed by 0.13 m, what is the spring constant of the spring?

OpenStudy (anonymous):

Use the formula PE=1/2(kx^2) 33.5J=1/2(k)(.13m)^2 Solve for k

OpenStudy (anonymous):

how do you solve for k? @Grazes

OpenStudy (anonymous):

... are you serious? Have you taken pre-algebra/algebra yet?